OpenStudy (yacoub1993):

Find a rational zero of the polynomial function and use it to find all the zeros of the function. f(x) = x4 + 3x3 - 5x2 - 9x - 2

OpenStudy (yacoub1993):

\[A. {1, -2, -2 +\sqrt{3} , -2 -\sqrt{3} }\] \[B. {-1, -2, -2 +\sqrt{5} , -2 -\sqrt{5} } \] \[C. {-1, 3, -2 +\sqrt{5} , -2 -\sqrt{5} }\] \[D. {-1, 2, -2 + \sqrt{3}, -2 -\sqrt{3} } \]

OpenStudy (yacoub1993):

@ozzman77 Help Please

OpenStudy (anonymous):

im sorry its been a long time since i have done these types of problems nd i wouldnt want 2 give you the wrong info. so id have 2 say i dnt know this 1

OpenStudy (yacoub1993):

ok thnxs

OpenStudy (yacoub1993):

@austinL Help

OpenStudy (yacoub1993):

@Easyaspi314

OpenStudy (austinl):

Okay.... I know a way that is EXTREMELY convoluted and rather difficult. It would take quite a while to explain as well....

OpenStudy (yacoub1993):

ok try to shorthen it while explaining

OpenStudy (austinl):

Unfortunately it can't really be shortened. I can however link you to someone else giving the method I was gonna use.